OpenStudy (lord2clash):

What is the equation of a line that is perpendicular to and passes through the point (8, 3)?

OpenStudy (lord2clash):

@TheSmartOne could i get help with this?

OpenStudy (steve816):

perpendicular to what?

OpenStudy (michele_laino):

please you have to provide the line of reference

OpenStudy (steve816):

Exactly

OpenStudy (lord2clash):

sorry didn't realize it wasn't there. its y=4x+5

OpenStudy (shadowlegendx):

Good ole algebra, miss those days

OpenStudy (michele_laino):

hint: if I call with \(m\) the slope of the refernce line, and with \(m'\) the slope to the perpendicular line to such reference line, then the subsequent condition holds: \(m \cdot m'=-1\)

OpenStudy (michele_laino):

reference*

OpenStudy (michele_laino):

now, I ask what is the slope \(m\) of the line of reference, please?

OpenStudy (lord2clash):

-1 since they are perpendicular they have the same slope

OpenStudy (michele_laino):

hint: what is the slope of this line: \(y=4x+5\) ?

OpenStudy (lord2clash):

4 (sorry i got confused xD)

OpenStudy (michele_laino):

that's right! :) from my formula above we get: \[m' = - \frac{1}{m} = ...?\] please replace \(m=4\), and compute the value of \(m'\). What do you get?

OpenStudy (lord2clash):

1/4 = 0.25

OpenStudy (michele_laino):

more precisely it is \(-1/4\), am I right?

OpenStudy (lord2clash):

yes

OpenStudy (michele_laino):

ok! Now in order to write the requested equation, we can apply this formula: \[y - y_0 = - \frac{1}{4}\left( {x - x_0} \right)\] where \(x_0=8,\;y_0=3\), please replace such values into such equation. What do you get?

OpenStudy (lord2clash):

y-3=-1/4 (x-8)

OpenStudy (michele_laino):

that's right! Now we have to simplify such equation, in order to get an equation of this type: \(y=mx+q\) so, I apply the distributive property of multiplication over addition to the right side, and I get: \[y - 3 = - \frac{x}{4} + 2\] now, please continue

OpenStudy (lord2clash):

I don't know what to do from here...

OpenStudy (michele_laino):

If I add \(3\) to both side, I can write this: \[y - 3 + 3 = - \frac{x}{4} + 2 + 3\] please simplify

OpenStudy (michele_laino):

sides*

OpenStudy (michele_laino):

hint: what is \(-3+3=...?\)

OpenStudy (shadowlegendx):

@Michele_Laino Youʻre a really good helper ^-^

OpenStudy (michele_laino):

thanks!! :) :) @ShadowLegendX